OpenStudy (anonymous):

Which of the following is the equation of the circle with center (4, 0) and r = sqrt 5? x^2 + (y - 4)^2 = 5 x^2 + (y + 4)^2 = sqrt 5 (x + 4)^2 + y^2 = sqrt 5 (x - 4)^2 + y^2 = 5

OpenStudy (phi):

the equation of a circle with center (h,k) and radius r is \[ (x-h)^2 + (y-k)^2 = r^2 \]

OpenStudy (jdoe0001):

\(\bf center \to ({\color{red}{ 4}},{\color{blue}{ 0}})\qquad radius=\sqrt{5} \\ \quad \\ (x-{\color{red}{ h}})^2+(y-{\color{blue}{ k}})^2=r^2\)

OpenStudy (phi):

replace h with 4, k with 0, and r with sqr(5)

OpenStudy (phi):

notice that sqr(5)^2 is sqr(5)*sqr(5) = 5 so the answer must be either the first or last choice

OpenStudy (anonymous):

Thank you!! Both of you. :0

OpenStudy (jdoe0001):

yw